OpenStudy (ayubie):

Calc 3 / sort of differential equations question Solve using a power series: I am given y′′ − (x^2 + 8)y′ + y = 0 centered around x0= 3 I've taken the derivatives and plugged them into the original ODE. Where I'm stuck is simplifying the second term (which is more of a Calc 3 question I think): -[(x2 + 8)] multiplied by the sum from zero to infinity of na(sub n) (x-3)^n-1

OpenStudy (rsadhvika):

Hint : \[x^2+8 = (x-3)^2 + 6(x-3)+17\]

OpenStudy (ayubie):

I'm assuming I need to solve for x? But what do I do once I get x? thanks :)

OpenStudy (rsadhvika):

Nope, just replace \(x^2+8\) by \((x-3)^2+6(x-3)+17\)

OpenStudy (ayubie):

ohhhhhh that makes sense. so I can plug in the (x-3)^2 and just change the value of n, what about the rest of that equation though?

OpenStudy (rsadhvika):

\(y = \sum\limits_{n=0}^{\infty}a_n(x-3)^n\) \(y' = \sum\limits_{n=1}^{\infty}a_n(n-1)(x-3)^{n-1}\) \(y'' = \sum\limits_{n=2}^{\infty}a_n(n-1)(n-2)(x-3)^{n-2}\)

OpenStudy (rsadhvika):

Plugging them in the given eqn gives : \[\sum\limits_{n=2}^{\infty}a_n(n-1)(n-2)(x-3)^{n-2}\\-(x^2+2) \sum\limits_{n=1}^{\infty}a_n(n-1)(x-3)^{n-1} \\+ \sum\limits_{n=1}^{\infty}a_n(x-3)^{n}\\=0\]

OpenStudy (rsadhvika):

Next replace \(x^2+8\) by \((x-3)^2+6(x-3)+17\)

OpenStudy (ayubie):

Wait so where did the \[-(x ^{2} + 2)\] come from? Or was that a typo and should have been an 8?

OpenStudy (rsadhvika):

sry its a typo, supposed to be x^2+8 : \[\sum\limits_{n=2}^{\infty}a_n(n-1)(n-2)(x-3)^{n-2}\\ -(x^2+\color{red}{8}) \sum\limits_{n=1}^{\infty}a_n(n-1)(x-3)^{n-1} \\ + \sum\limits_{n=1}^{\infty}a_n(x-3)^{n}\\ =0\]

OpenStudy (ayubie):

okay cool that makes sense, so then for the middle term I understand how to move (x-3)^2 inside of the series, how do I move the rest inside?

OpenStudy (rsadhvika):

\[\sum\limits_{n=2}^{\infty}a_n(n-1)(n-2)(x-3)^{n-2}\\ -[(x-3)^2+6(x-3)+17] \sum\limits_{n=1}^{\infty}a_n(n-1)(x-3)^{n-1} \\ + \sum\limits_{n=1}^{\infty}a_n(x-3)^{n}\\ =0\]

OpenStudy (rsadhvika):

\[\sum\limits_{n=2}^{\infty}a_n(n-1)(n-2)(x-3)^{n-2}\\ - \sum\limits_{n=1}^{\infty}a_n(n-1)(x-3)^{n+1} -6 \sum\limits_{n=1}^{\infty}a_n(n-1)(x-3)^{n}-17 \sum\limits_{n=1}^{\infty}a_n(n-1)(x-3)^{n-1}\\ + \sum\limits_{n=1}^{\infty}a_n(x-3)^{n}\\ =0\]

OpenStudy (rsadhvika):

Notice that you have 5 series in above eqn

OpenStudy (rsadhvika):

They all have different exponents : \((x-3)^{n-2},~~(x-3)^{n+1},~~(x-3)^n,~~(x-3)^{n-1}\)

OpenStudy (rsadhvika):

Your next step is to make them all have same exponent : \((x-3)^n\)

OpenStudy (rsadhvika):

Do you know how to do that ?

OpenStudy (ayubie):

Is that where I change the lower boundary?

OpenStudy (rsadhvika):

Yes

OpenStudy (rsadhvika):

simply use substitution

OpenStudy (rsadhvika):

Let me show you how to do it for first series

OpenStudy (ayubie):

So I also just realized I had the derivatives written different, I had \[y''= \sum_{0}^{\infty} (n-1)na _{n}(x-3)^{n-2}\], is this wrong?

OpenStudy (rsadhvika):

Hey wait, there is a mistake in my derivatives

OpenStudy (rsadhvika):

Haven't you noticed my mistake yet ?

OpenStudy (ayubie):

I think I noticed it above, right?

OpenStudy (rsadhvika):



OpenStudy (rsadhvika):

see the mistake ?

OpenStudy (ayubie):

You're missing the n?

OpenStudy (rsadhvika):

Yes

OpenStudy (rsadhvika):

Lets fix the differential equation quick

OpenStudy (ayubie):

Are there two different ways to do the derivatives? Because I wrote mine differently before

OpenStudy (rsadhvika):

Here is the corrected equation : \[\sum\limits_{n=2}^{\infty}a_n n(n-1)(x-3)^{n-2}\\ - \sum\limits_{n=1}^{\infty}a_n n(n-1)(x-3)^{n+1} -6 \sum\limits_{n=1}^{\infty}a_n n(n-1)(x-3)^{n}-17 \sum\limits_{n=1}^{\infty}a_n n(n-1)(x-3)^{n-1}\\ + \sum\limits_{n=1}^{\infty}a_n(x-3)^{n}\\ =0\]

OpenStudy (ayubie):

Okay that makes sense

OpenStudy (rsadhvika):

Nope, there is only one way to do differentiation. See if above looks good

OpenStudy (rsadhvika):

Now try making all the exponents equal to (x-3)^n

OpenStudy (ayubie):

So I would change the lower bound for the first term to negative two to get the exponent the same?

OpenStudy (rsadhvika):

I'll show you how to fix the exponent for the first series

OpenStudy (rsadhvika):

\(\sum\limits_{n=2}^{\infty}a_n n(n-1)(x-3)^{n-2}\) Let \(k=n-2\) this gives \(n = k+2\) and the lower bound becomes \(k=2-2=0\) the series is then : \(\sum\limits_{k=0}^{\infty}a_{k+2} (k+2)(k+1)(x-3)^{k}\)

OpenStudy (rsadhvika):

Btw, index variable is just dummy, it doesn't matter what variable you use for index. So you can replace k by n in the end

OpenStudy (rsadhvika):

\(\sum\limits_{k=0}^{\infty}a_{k+2} (k+2)(k+1)(x-3)^{k}\) is same as \(\sum\limits_{n=0}^{\infty}a_{n+2} (n+2)(n+1)(x-3)^{n}\)

OpenStudy (rsadhvika):

try fixing the exponents of remaining series

OpenStudy (ayubie):

okay gotcha

OpenStudy (ayubie):

So would the next part be \[\sum_{2}^{\infty} a_{n-1}(n-1)(n-1)(x-3)^n\]

OpenStudy (ayubie):

wait, but the second (n-1) should be (n-2) my bad

OpenStudy (rsadhvika):

Lets see

OpenStudy (rsadhvika):

\(\sum\limits_{n=1}^{\infty}a_n n(n-1)(x-3)^{n+1}\) Let \(n+1=k\) this gives \(n=k-1\) and the lower bound becomes \(k=1+1=2\) the series is then \(\sum\limits_{k=2}^{\infty}a_{k-1} (k-1)(k-2)(x-3)^{k}\)

OpenStudy (rsadhvika):

that is same as \(\sum\limits_{n=2}^{\infty}a_{n-1} (n-1)(n-2)(x-3)^{n}\)

OpenStudy (ayubie):

okay awesome, I think I got all of them then. So now I need to change the lower bound to n=2 because that's the highest one, correct?

OpenStudy (rsadhvika):

Exactly!

OpenStudy (ayubie):

which is just pulling out factors

OpenStudy (rsadhvika):

To do that, just observe this split up : \[\sum\limits_{n=0}^{\infty}a_n x^n = a_0+a_1x + \sum\limits_{n=2}^{\infty}a_n x^n \]

OpenStudy (rsadhvika):

Yes its just pulling out first few terms

OpenStudy (ayubie):

okay sweet, ill try the first one then

OpenStudy (ayubie):

Okay so I get \[2a_{2} + 6a_{3}(x-3) + \sum_{n=2}^{\infty} a_{n+2}(n+2)(n+1)(x-3)^n\] for the first term

OpenStudy (ayubie):

And then for the second term I don't do anything, and for the third term plugging in 1 for n will give me zero?

OpenStudy (rsadhvika):

Looks perfect. Yes, for the third term : \[\sum\limits_{n=1}^{\infty}a_n n(n-1)(x-3)^{n}= 0+ \sum\limits_{n=2}^{\infty}a_n n(n-1)(x-3)^{n}\]

OpenStudy (rsadhvika):

Hey sorry wait

OpenStudy (ayubie):

So the bottom part is what I have right now, is that correct?

OpenStudy (rsadhvika):

scratch all that, there is a mistake in our derivative for y' itself

OpenStudy (ayubie):

oh but the x^n should be (x-3)^n at the end

OpenStudy (ayubie):

ohhhhh

OpenStudy (rsadhvika):

Wrong equation : \[\sum\limits_{n=2}^{\infty}a_n n(n-1)(x-3)^{n-2}\\ - \sum\limits_{n=1}^{\infty}a_n n\color{red}{(n-1)}(x-3)^{n+1} -6 \sum\limits_{n=1}^{\infty}a_n n\color{red}{(n-1)}(x-3)^{n}-17 \sum\limits_{n=1}^{\infty}a_n n\color{red}{(n-1)}(x-3)^{n-1}\\ + \sum\limits_{n=1}^{\infty}a_n(x-3)^{n}\\ =0\] Correct equation : \[\sum\limits_{n=2}^{\infty}a_n n(n-1)(x-3)^{n-2}\\ - \sum\limits_{n=1}^{\infty}a_n n(x-3)^{n+1} -6 \sum\limits_{n=1}^{\infty}a_n n(x-3)^{n}-17 \sum\limits_{n=1}^{\infty}a_n n(x-3)^{n-1}\\ + \sum\limits_{n=1}^{\infty}a_n(x-3)^{n}\\ =0\]

OpenStudy (rsadhvika):

you will need to work from above equation again, sry about that :)

OpenStudy (ayubie):

no problem. so from there just change the exponents again?

OpenStudy (rsadhvika):

Yes

OpenStudy (rsadhvika):

1) change the exponent of all series to (x-3)^n 2) have same lower bound for all series by pulling out first few terms

OpenStudy (ayubie):

okay ill do that now

OpenStudy (rsadhvika):

Found another mistake

OpenStudy (rsadhvika):

Here is the final eqn with all mistakes corrected, hopefully : \[ \sum\limits_{n=2}^{\infty}a_n n(n-1)(x-3)^{n-2}\\ - \sum\limits_{n=1}^{\infty}a_n n(x-3)^{n+1} -6 \sum\limits_{n=1}^{\infty}a_n n(x-3)^{n}-17 \sum\limits_{n=1}^{\infty}a_n n(x-3)^{n-1}\\ + \sum\limits_{n=\color{red}{0}}^{\infty}a_n(x-3)^{n}\\ =0\]

OpenStudy (ayubie):

alright sounds good, I hadn't reached there yet anyways haha

OpenStudy (ayubie):

\[2a_{2}+a_{0}+6a_{3}(x-3)-5a_{1}(x-3)+\sum_{n=2}^{\infty}[a_{n+2}(n+2)(n+1)-a_{n-1}(n-1)-6a_{n}n-17a_{n+1}(n+1)+a_{n}](x-3)^n\] This is what I got from that

OpenStudy (rsadhvika):

Looks you have forgot the pulled out terms from fourth series : \[-17a_1-34a_2(x-3)\]

OpenStudy (ayubie):

oh right I wrote it down just forgot it in the equation writer haha

OpenStudy (rsadhvika):

This problem is so lenghty, but I see you have the right ideas... I hope they don't give you this in exam

OpenStudy (ayubie):

so now I can set the entire inside of that sum equal to zero and that will be my RR for n=2,3,...

OpenStudy (ayubie):

I dont think they will, i hope not

OpenStudy (rsadhvika):

Yes, before that set the coeffecients of (x-3)^0, (x-3)^1 equal to 0 and solve whatever coefficents you can

OpenStudy (ayubie):

and then I can set a0=c1 and a1=c2 because it's a second order equation?

OpenStudy (rsadhvika):

Yes

OpenStudy (rsadhvika):

setting the coefficient of (x-3)^0 equal to 0 : \(2a_2+a_0-17a_1 = 0\) \(\implies a_2 = \dfrac{-c_1+17c_2}{2}\)

OpenStudy (rsadhvika):

setting the coefficient of (x-3)^1 equal to 0 : \(6a_3-5a_1-34a_2 = 0\) \(\implies a_3=\cdots \)

OpenStudy (ayubie):

okay sweet, and I just need the first 3 terms of a for the answer

OpenStudy (ayubie):

\[a_{3}=[5c_{2}-17(17c_{2}-c_{1})]/6\] would be the third term then?

OpenStudy (ayubie):

it says I need the first 3 terms in each of two linearly independent solutions

OpenStudy (rsadhvika):

For the first solution, set \(c_1 = 1, ~c_2=0\) For the second solution, set \(c_1 = 0, ~c_2=1\)

OpenStudy (ayubie):

so just plug in those values and my answer will be the results?